# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v4

@@ 1 Introduction
Deep learning-based algorithms have achieved remarkable success across various disciplines, particularly in computer vision (CV) [1] and natural language processing (NLP) [2]. These models often utilize a pretraining task on large datasets to improve performance. This initial step typically serves as an initialization point, after which the model is fine-tuned for specific downstream tasks. Among pre-training strategies for deep learning, self-supervised learning (SSL) has emerged as a prominent approach. SSL approach is driven by two key motivations. Firstly, network trained on extensive datasets learn transferable patterns that reduce overfitting during training. Secondly, parameters derived from large-scale datasets provide effective initialization, enabling faster convergence between different applications [3].
Despite the abundance of unlabeled web data in the era of big data, developing high-quality labeled datasets through human annotation remains expensive. For instance, platforms such as Scale.com [4] charge approximately \$6.4 per image for the labeling task such as image segmentation. Building an image segmentation dataset containing over a million high-quality samples, like JFT-300M, can cost millions of dollars. Usually, this process is time-intensive and inefficient [5]. Moreover, methods of supervised learning can easily learn misleading connections within dataset, which can result in mistakes and thus may be more sensitive to adversarial exploitations.
To address these challenges, strategies like active learning [6], semi-supervised learning [7], and SSL [8] have gained traction. Recently, transformers [9] have emerged as an effective neural network (NN) architecture to leverage deeper insights into input data; specifically, ViTs [10] are considered the alternative to CNNs [11]. ViTs have been successfully employed in tasks related to computer vision, including object detection [12], recognition [13] and semantic segmentation [14]. In particular, ViTs have demonstrated superior performance in image classification, when trained with large-volume datasets such as JFT-300M [15].
In essence, SSL methods enable Vision Transformers (ViTs) to understand the underlying structures of data without supervision through pretext tasks, the self-designed objectives in a pre-training phase that encourage the model to learn significant patterns within the data itself. The motivation for applying SSL in ViTs is two-fold. Firstly, SSL enables ViTs to utilize large amounts of unlabeled data, fostering robust and generalizable feature representations. This approach is particularly advantageous when labeled data is limited or costly to produce. Secondly, SSL facilitates pre-training ViTs on extensive datasets, which can then be fine-tuned for specific downstream tasks. This process not only improves the performance of the model but also accelerates training convergence [5].
1.1 Survey Scope and Contributions
@@
Several interesting surveys related to SSL have been reported. However, these surveys on SSL, serve to particular domains like recommendation systems [16], graphs [17], [18], sequential transfer learning [19], videos [20] and algorithms [5]. Vision Transformers (ViTs) are a prominent area of research in computer vision and have recently demonstrated several breakthroughs. However, the training of ViTs remains a challenge. Therefore, in contrast to existing surveys, this paper explores the different SSL methods proposed for ViTs.
@@ 1.2 Survey Structure
The paper begins with an introduction to SSL and a detailed discussion on importance of SSL in ViTs. Then we discussed the recent advancements and applications of SSL across different computer vision tasks. It also presents a taxonomy for SSL, which classifies them according to how feature learning is applied within their architecture. According to this taxonomy, SSLs are divided into five groups, each representing unique way to make use of input features. Frequently used abbreviations are listed in Table. 1. The structure of paper is depicted in Fig. LABEL:fig:intro. Section 1 discuss a systematic understanding of the SSL architecture, highlighting its need in ViTs and outlined the advent of SSL architectures. Proceed to section 2 which covers the advancements in SSL variants, while section 3 provides a taxonomy of the recent SSL architectures, respectively.
Section 4 focuses on the regularization techniques used in SSL, particularly in the area of computer vision. Section 5 presents various metrics and benchmarks to evaluate the effectiveness of SSL, while Section 6 provides a comparative analysis of state-of-the-art SSL mechanisms that employ Vision Transformers. Section 7 highlights some open challenges in the domain of SSL, and Section 8 discusses different applications and future directions for SSL in Vision Transformers, including current challenges and potential developments. Finally, Section 9 concludes this survey.
@@
BERT Bi-Directional Representation of Transformers
@@
MCVT Multi-level Contrastive Learning
@@
DINO Slef-Distillation with NO Labels
@@
1-2 GCMAE Global Contrast Masked Autoencode
HSL Hybrid SSL Framewor
@@ 2 Evolution of SSL
In computer vision SSL methods are generally segmented into contrastive, generative, and predictive approaches. Contrastive methods, such as MoCo [21] and SimCLR [22], aim to learn patterns by contrasting positive and negative samples. Generative methods, such as GANs [23] and VAEs [24], focus on learning representations by generating samples from the learned representation space. Predictive methods, such as BERT [25] and T5 [26], learn representations by predicting some parts of the input data.
@@ 2.1 Evolution of ViTs
In the domain of computer vision, ViTs is newly arising development motivated by the success of transformers architecture which is presented in seminal paper ”Attention Is All You Need” [9]. Transformers has revolutionized the learning process by enabling models to learn patterns from substantial volume of unlabeled data.
In computer vision, the traditional approach in image processing tasks is the utilization of convolutional neural networks (CNNs). However, the CNNs have limitations in capturing dependencies across long spatial distances and global context, both of which are crucial for some vision tasks. ViTs address these limitations via self-attention mechanisms to learn and understand the global context and also dependencies over long spatial distances in the input data .
@@ 2.2 Importance of SSL in ViTs
SSL is especially crucial in the context of ViTs, as it empowers models to learn and understand feature representations from extensive unlabeled data. ViTs rely on extensive datasets to learn meaningful and effective representations while SSL offers a solution by enabling pretraining on large-scale datasets, allowing ViTs to acquire robust and transferable feature representations for downstream tasks.
@@ 2.3 Evolution of SSL in Computer Vision
Jigsaw Puzzles Object parts, Spatial layout Temporal dynamics
Patch Position Prediction Spatial hierarchies Color/texture learning
Colorization Textures, Patterns, Object ID Spatial relationships
Inpainting Texture continuity, Object completion Temporal understanding
Frame Prediction Motion patterns, Temporal consistency Static image features
Temporal Order Verification Temporal understanding Static spatial understanding
SimCLR, MoCo, CPC Feature discrimination Temporal, Cross-modal learning
DeepCluster, SwAV Feature grouping, Discrimination Dynamic scene understanding
BYOL, DINO, SimSiam Consistent feature learning High-frequency detail understanding
Barlow Twins, VICReg Unique feature emphasis Temporal feature learning
Audio-Visual Correspondence Cross-modal learning Isolated visual feature learning
Self-Labelling Simultaneous clustering and representation Limited scalability
Local Aggregation Localized feature learning Global contextual understanding
TiCo Transformation invariance, Covariance contrast Isolated feature learning
Contrastive Multiview Coding Multiview feature learning Single modality data processing
Relative Predictive Coding Relative feature prediction Absolute feature understanding
@@ 2.4 Comparison of SSL with Transfer Learning
Challenges associated with TL or SSL especially in complex vision problems are due to several factors including the volume of pre-training data and class imbalances in the source task. A study by [58] examined this issue across multiple sources and target tasks, yielding the following key insights:
@@ 3.1 Contrastive Methods
Contrastive Learning (CL) stands as a prominent technique in SSL [64, 21, 22]. Its aim is to capture invariant semantics through pairs of random views, termed Contrastive multi-view coding. CL ensures that global projections of representations are similar for positive samples and dissimilar for negative samples. This section delves into literature reviews on Contrastive Learning applied in transformer frameworks.
Literature based Review of Contrastive Learning Li et al. [65] suggested a methodology to enhance domain generalization in mammographic image analysis using contrastive learning. Their approach consists of two stages: first, a multi-style and multi-view contrastive learning pretrainig phase, where CycleGAN and image blending techniques generate diverse mammogram styles to enhance feature robustness; second, a fine-tuning stage for downstream tasks such as mass detection, mass matching, BI-RADS classification, and breast density classification. By learning style-variant representations, their method demonstrated improved performance across multiple mammographic datasets, outperforming existing domain generalization techniques. Wang and co-workers in 2022 [66] introduced the method SRCL that is semantically-relevant contrastive learning for histopathology image analysis. Unlike traditional contrastive learning, SRCL evaluates similarity among instances to identify additional positive pairs by arranging various positive examples that share related visual concepts. This approach enriches the diversity of positive pairs and yields more informative patterns. They employee a hybrid model named CTransPath, in which they combines a CNN with a multi-scale Swin Transformer architecture, their methodology acts as a unified feature extractor, addressing both local and global features during pre-training, seeking to acquire global feature interpretations tailored for activities within the domain of histopathology images. The efficacy of SRCL-pretrained CTransPath was evaluated across five distinct downstream tasks spanning nine publicly available datasets, such as retrieving patches, classifying patches, performing weakly-supervised classification of whole-slide images, detecting mitosis, and segmenting glands in colorectal adenocarcinoma. The findings demonstrated that the visual representations derived from SRCL not just achieved state-of-the-art performance throughout all datasets but it also exhibited increased robustness and transferability compared to both supervised and self-supervised ImageNet pre-training techniques.
@@ 3.1.1 Knowledge Distillation Methods
Knowledge distillation transfers knowledge from a complex teacher model to a simpler student model without labeled data. This section explores literature on knowledge distillation in SSL applied to ViTs. Literature based Review of Knowledge Distillation Caron et al. [45] in 2021 introduced knowledge distillation into SSL to enhance the training of ViTs. Their method, DINO, comprises two identical networks termed student and teacher networks, each with distinct parameters. Both networks include an encoder or base network, utilizing either ViTs or ResNet50, along with a projection head placed on top of the encoder network. The DINO approach involves generating multiple crops from two perspectives: a local perspective and a global perspective. The global crop perspective is fed to the teacher network, while the local crop perspective is fed to the student network. The local crop is a subset of the global perspective, exhibiting overlap between the two. The teacher network guides updates in the student network, but gradient stop is applied to halt backpropagation in the teacher network. Meanwhile, the teacher network weights are adjusted using the EMA (exponential moving average) of the weights from student network.
Li et al. (2022) [12] proposed Efficient Self-Supervised ViTs (EsViT) to achieve a optimized balance between accuracy and efficiency, reducing the cost in building state-of-the-art SSL vision systems while demonstrating enhanced scaling performance on accuracy versus throughput and model size. They introduced a multi-stage Transformer architecture, combining networks to address the issue of semantic information loss when transitioning from a single-stage transformer to a multi-stage architecture. Employing a distillation learning strategy targeting correspondence between patches, they introduced a new non-contrastive region-matching pre-training task, enhancing the model’s ability to learn intricate regional dependencies and substantially improve the quality of visual representations learned.
The Multi-stage ViT segments an input RGB image into individual non-overlapping patches, handling each patch as an individual token in the first stage. Later, the patch merging module concatenates the features of each group and applies a linear layer to reduce the number of tokens. Subsequently, a Transformer with a sparse self-attention module enables interactions among the merged features. This process is repeated multiple times, resulting in a hierarchical representation. EsViT, the proposed model, is validated across various tasks. EsViT achieves 81.3% top-1 accuracy, with the linear evaluation protocol, demonstrating 3.5× parameter efficiency and at least 10× higher throughput compared to previous state-of-the-art models. EsViT leads over its supervised counterpart, Swin Transformers, by outperforming on 17 of 18 datasets in downstream linear classification tasks.
@@ 3.1.2 Clustering Methods
Su et al. (2023) [73] introduced a clustering-based SSL framework tailored for ViTs, aiming to improve the performance of tasks related to object detection and segmentation. Their proposed methodology, Feature-Level SSL (FLSL), works at two levels of semantics: intra-view, which deals with individual image, and inter-view, which considers over an entire dataset. FLSL utilizes a bilevel feature clustering approach, integrating mean-shift clustering intrinsic to transformers for extracting modes as patterns with a k-means-based SSL technique, ensuring semantic coherence of extracted representations both locally and globally.
At the first semantic level, FLSL optimizes intra-cluster affinity using a self-attention layer to encourage semantic representations within clusters. On the second tier, semantic representations are cultivated through non-empty k-means clustering, with positive samples identified using a cross-attention layer. Experimental outcomes illustrate the efficacy of FLSL in dense prediction tasks, achieving significant improvements in object detection and instance-level segmentation on MS-COCO, utilizing Mask R-CNN with ViT-S/16 and ViT-S/8 as backbones, respectively.

### v5

@@ 1 Introduction
Deep learning-based algorithms have achieved remarkable success across various disciplines, particularly in computer vision (CV) [1, 2] and natural language processing (NLP) [3]. These models often utilize a pretraining task on large datasets to improve performance. This initial step typically serves as an initialization point, after which the model is fine-tuned for specific downstream tasks. Among pre-training strategies for deep learning, self-supervised learning (SSL) has emerged as a prominent approach. SSL approach is driven by two key motivations. Firstly, network trained on extensive datasets learn transferable patterns that reduce overfitting during training. Secondly, parameters derived from large-scale datasets provide effective initialization, allowing faster convergence between different applications [4]. Despite the abundance of unlabeled web data in the era of big data, developing high-quality labeled datasets through human annotation remains expensive. For instance, platforms such as Scale.com [5] charge approximately \$6.4 per image for the labeling task such as image segmentation. Building an image segmentation dataset containing over a million high-quality samples, like JFT-300M, can cost millions of dollars. Usually, this process is time-intensive and inefficient [6]. Moreover, methods of supervised learning can easily learn misleading connections within dataset, which can result in mistakes and thus may be more sensitive to adversarial exploitations. To address these challenges, strategies like active learning [7], semi-supervised learning [8], and SSL [9] have gained traction. Recently, transformers [10] have emerged as an effective neural network (NN) architecture to leverage deeper insights into input data; specifically, ViTs [11] are considered the alternative to CNNs [12, 13, 14, 15, 16, 17] ViTs have been successfully employed in tasks related to computer vision, including object detection [18], recognition [19] and semantic segmentation [20]. In particular, ViTs have demonstrated superior performance in image classification, when trained with large-volume datasets such as JFT-300M [21]. In essence, SSL methods enable Vision Transformers (ViTs) to understand the underlying structures of data without supervision through pretext tasks, the self-designed objectives in a pre-training phase that encourage the model to learn significant patterns within the data itself. The motivation for applying SSL in ViTs is two-fold. Firstly, SSL enables ViTs to utilize large amounts of unlabeled data, fostering robust and generalizable feature representations. This approach is particularly advantageous when labeled data is limited or costly to produce. Secondly, SSL facilitates pre-training ViTs on extensive datasets, which can then be fine-tuned for specific downstream tasks. This process not only improves the performance of the model but also accelerates training convergence [6].
@@
Several interesting surveys related to SSL have been reported. However, these surveys on SSL, serve to particular domains like recommendation systems [22], graphs [23], [24], sequential transfer learning [25], videos [26], algorithms [6], non-sequential tabular data [27] and recent advancements in using deep learning and image-based models for plant disease detection [28]. A recent survey summarize the deep learning models deployed on edge devices, particularly in healthcare and computer vision [29], highlighting the trade-offs between performance and resources usage. Vision Transformers (ViTs) are a prominent area of research in computer vision [30] and have recently demonstrated several breakthroughs. However, the training of ViTs remains a challenge.
In contrast to existing surveys, this paper presents a comprihensive examination of different SSL methods proposed for ViTs, offering a structured taxonomy for pre-text techniques, a comparitive analysis with transfer learning paradigms, an overview of regularization strategies and evaluation metrics, and a discussion on prevailing challenges along with prospective research directions.
Figure 1: Structure of various sections of the survey paper.
@@ 1.1 Survey Structure
The paper begins with an introduction to SSL and a detailed discussion on importance of SSL in ViTs. Then we discussed the recent advancements and applications of SSL across different computer vision tasks. It also presents a taxonomy for SSL, which classifies them according to how feature learning is applied within their architecture. According to this taxonomy, SSLs are divided into five groups, each representing unique way to make use of input features. Frequently used abbreviations are listed in Table. 1. The structure of paper is depicted in Fig. 1. Section 1 discuss a systematic understanding of the SSL architecture, highlighting its need in ViTs and outlined the advent of SSL architectures. Proceed to section 2 which covers the advancements in SSL variants, while section 3 provides a taxonomy of the recent SSL architectures, respectively. Section 4 focuses on the regularization techniques used in SSL, particularly in the area of computer vision. Section 5 presents various metrics and benchmarks to evaluate the effectiveness of SSL, while Section 6 provides a comparative analysis of state-of-the-art SSL mechanisms that employ Vision Transformers. Section 7 highlights some open challenges in the domain of SSL, and Section 8 discusses different applications and future directions for SSL in Vision Transformers, including current challenges and potential developments. Finally, Section 9 concludes this survey.
@@
BERT Bi-Directional Encoder Representations from Transformers
@@
MCVT Multi-level Contrastive Learning for Vision Transformers
@@
DINO Self-Distillation with NO Labels
@@
GCMAE Global Contrast Masked Autoencoder
HSL Hybrid SSL Framework
@@ 2 Evolution of SSL
In computer vision, SSL methods are generally segmented into contrastive, generative, and predictive approaches. Contrastive methods, such as MoCo [31] and SimCLR [32], aim to learn patterns by contrasting positive and negative samples. Generative methods, such as GANs [33] and VAEs [34], focus on learning representations by generating samples from the learned representation space. Predictive methods, such as BERT [35] and T5 [36], learn representations by predicting some parts of the input data.
@@ 2.1 Evolution of ViTs
In the domain of computer vision, ViTs is newly arising development motivated by the success of transformers architecture which is presented in seminal paper ”Attention Is All You Need” [10]. Transformers has revolutionized the learning process by enabling models to learn patterns from substantial volume of unlabeled data. This approach enables global context modeling, distinguishing ViTs from traditional CNNs. Early ViT models demonstrated strong performance in image classification but faced challenges in computational efficiency and data requirements [39].
In computer vision, the traditional approach in image processing tasks is the utilization of convolutional neural networks (CNNs). However, the CNNs have limitations in capturing dependencies across long spatial distances and global context, both of which are crucial for some vision tasks. ViTs address these limitations via self-attention mechanisms to learn and understand the global context and also dependencies over long spatial distances in the input data. ViTs have also shown strong results in image classification, object detection, medical imaging, and image restoration, often surpassing CNNs when trained on large datasets [40].
@@ 2.2 Importance of SSL in ViTs
SSL is especially crucial in the context of ViTs, as it empowers models to learn and understand feature representations from extensive unlabeled data. ViTs rely on extensive datasets to learn meaningful and effective representations while SSL offers a solution by enabling pretraining on large-scale datasets, allowing ViTs to acquire robust and transferable feature representations for downstream tasks [44, 45, 46, 47].
@@ 2.3 Evolution of SSL in Computer Vision
Jigsaw Puzzles [53] Object parts, Spatial layout Temporal dynamics
Patch Position Prediction [37] Spatial hierarchies Color/texture learning
Colorization [52] Textures, Patterns, Object ID Spatial relationships
Inpainting [51] Texture continuity, Object completion Temporal understanding
Frame Prediction [72] Motion patterns, Temporal consistency Static image features
Temporal Order Verification [73] Temporal understanding Static spatial understanding
SimCLR, MoCo, CPC [32, 31, 60] Feature discrimination Temporal, Cross-modal learning
DeepCluster, SwAV [55, 37] Feature grouping, Discrimination Dynamic scene understanding
BYOL, DINO, SimSiam [38, 61, 74] Consistent feature learning High-frequency detail understanding
Barlow Twins, VICReg [65, 64] Unique feature emphasis Temporal feature learning
Audio-Visual Correspondence [75] Cross-modal learning Isolated visual feature learning
Self-Labelling [76] Simultaneous clustering and representation Limited scalability
Local Aggregation [77] Localized feature learning Global contextual understanding
TiCo [78] Transformation invariance, Covariance contrast Isolated feature learning
Contrastive Multiview Coding [79] Multiview feature learning Single modality data processing
Relative Predictive Coding [80] Relative feature prediction Absolute feature understanding
@@ 2.4 Comparison of SSL with Transfer Learning
The challenges associated with TL or SSL especially in complex vision problems are due to several factors, including the volume of pre-training data and class imbalances in the source task. A study by [83] examined this issue across multiple sources and target tasks, yielding the following key insights:
@@ 3.1 Contrastive Methods
Contrastive Learning (CL) stands as a prominent technique in SSL [89, 31, 32]. Its aim is to capture invariant semantics through pairs of random views, termed Contrastive multi-view coding. CL ensures that global projections of representations are similar for positive samples and dissimilar for negative samples. This section delves into literature reviews on Contrastive Learning applied in transformer frameworks [90].
Literature based Review of Contrastive Learning
Li et al. [91] suggested a methodology to enhance domain generalization in mammographic image analysis using contrastive learning. Their approach consists of two stages: first, a multi-style and multi-view contrastive learning pretrainig phase, where CycleGAN and image blending techniques generate diverse mammogram styles to enhance feature robustness; second, a fine-tuning stage for downstream tasks such as mass detection, mass matching, BI-RADS classification, and breast density classification. By learning style-variant representations, their method demonstrated improved performance across multiple mammographic datasets, outperforming existing domain generalization techniques. Wang and co-workers in 2022 [92] introduced the method SRCL that is semantically-relevant contrastive learning for histopathology image analysis. Unlike traditional contrastive learning, SRCL evaluates similarity among instances to identify additional positive pairs by arranging various positive examples that share related visual concepts. This approach enriches the diversity of positive pairs and yields more informative patterns. They employee a hybrid model named CTransPath, in which they combines a CNN with a multi-scale Swin Transformer architecture, their methodology acts as a unified feature extractor, addressing both local and global features during pre-training, seeking to acquire global feature interpretations tailored for activities within the domain of histopathology images. The efficacy of SRCL-pretrained CTransPath was evaluated across five distinct downstream tasks spanning nine publicly available datasets, such as retrieving patches, classifying patches, performing weakly-supervised classification of whole-slide images, detecting mitosis, and segmenting glands in colorectal adenocarcinoma. The findings demonstrated that the visual representations derived from SRCL not just achieved state-of-the-art performance throughout all datasets but it also exhibited increased robustness and transferability compared to both supervised and self-supervised ImageNet pre-training techniques.
@@ 3.1.1 Knowledge Distillation Methods
Knowledge distillation transfers knowledge from a complex teacher model to a simpler student model without labeled data. This section explores literature on knowledge distillation in SSL applied to ViTs.
Literature based Review of Knowledge Distillation Caron et al. [61] in 2021 introduced knowledge distillation into SSL to enhance the training of ViTs. Their method, DINO, comprises two identical networks termed student and teacher networks, each with distinct parameters. Both networks include an encoder or base network, utilizing either ViTs or ResNet50, along with a projection head placed on top of the encoder network. The DINO approach involves generating multiple crops from two perspectives: a local perspective and a global perspective. The global crop perspective is fed to the teacher network, while the local crop perspective is fed to the student network. The local crop is a subset of the global perspective, exhibiting overlap between the two. The teacher network guides updates in the student network, but gradient stop is applied to halt backpropagation in the teacher network. Meanwhile, the teacher network weights are adjusted using the EMA (exponential moving average) of the weights from student network.
Li et al. (2022) [18] proposed Efficient Self-Supervised ViTs (EsViT) to achieve a optimized balance between accuracy and efficiency, reducing the cost in building state-of-the-art SSL vision systems while demonstrating enhanced scaling performance on accuracy versus throughput and model size. They introduced a multi-stage Transformer architecture, combining networks to address the issue of semantic information loss when transitioning from a single-stage transformer to a multi-stage architecture. Employing a distillation learning strategy targeting correspondence between patches, they introduced a new non-contrastive region-matching pre-training task, enhancing the model’s ability to learn intricate regional dependencies and substantially improve the quality of visual representations learned. The Multi-stage ViT segments an input RGB image into individual non-overlapping patches, handling each patch as an individual token in the first stage. Later, the patch merging module concatenates the features of each group and applies a linear layer to reduce the number of tokens. Subsequently, a Transformer with a sparse self-attention module enables interactions among the merged features. This process is repeated multiple times, resulting in a hierarchical representation. EsViT, the proposed model, is validated across various tasks. EsViT achieves 81.3% top-1 accuracy, with the linear evaluation protocol, demonstrating 3.5× parameter efficiency and at least 10× higher throughput compared to previous state-of-the-art models. EsViT leads over its supervised counterpart, Swin Transformers, by outperforming on 17 of 18 datasets in downstream linear classification tasks.
@@ 3.1.2 Clustering Methods
Su et al. (2023) [98] introduced a clustering-based SSL framework tailored for ViTs, aiming to improve the performance of tasks related to object detection and segmentation. Their proposed methodology, Feature-Level SSL (FLSL), works at two levels of semantics: intra-view, which deals with individual image, and inter-view, which considers over an entire dataset. FLSL utilizes a bilevel feature clustering approach, integrating mean-shift clustering intrinsic to transformers for extracting modes as patterns with a k-means-based SSL technique, ensuring semantic coherence of extracted representations both locally and globally. At the first semantic level, FLSL optimizes intra-cluster affinity using a self-attention layer to encourage semantic representations within clusters. On the second tier, semantic representations are cultivated through non-empty k-means clustering, with positive samples identified using a cross-attention layer. Experimental outcomes illustrate the efficacy of FLSL in dense prediction tasks, achieving significant improvements in object detection and instance-level segmentation on MS-COCO, utilizing Mask R-CNN with ViT-S/16 and ViT-S/8 as backbones, respectively.